# Learning the Covariance of Treatment Effects Across Many Weak Experiments

### v1

@@ Abstract
When primary objectives are insensitive or delayed, experimenters may instead focus on proxy metrics derived from secondary outcomes. For example, technology companies often infer long-term impacts of product interventions from their effects on weighted indices of short-term user engagement signals. We consider meta-analysis of many historical experiments to learn the covariance of treatment effects on different outcomes, which can support the construction of such proxies. Even when experiments are plentiful and large, if treatment effects are weak, the sample covariance of estimated treatment effects across experiments can be highly biased and remains inconsistent even as more experiments are considered. We overcome this by using techniques inspired by weak instrumental variable analysis, which we show can reliably estimate parameters of interest, even without a structural model. We show the Limited Information Maximum Likelihood (LIML) estimator learns a parameter that is equivalent to fitting total least squares to a transformation of the scatterplot of estimated treatment effects, and that Jackknife Instrumental Variables Estimation (JIVE) learns another parameter that can be computed from the average of Jackknifed covariance matrices across experiments. We also present a total-covariance-based estimator for the latter estimand under homoskedasticity, which we show is equivalent to a $k$-class estimator. We show how these parameters relate to causal quantities and can be used to construct unbiased proxy metrics under a structural model with both direct and indirect effects subject to the INstrument Strength Independent of Direct Effect (INSIDE) assumption of Mendelian randomization. Lastly, we discuss the application of our methods at Netflix.
Keywords: surrogate outcomes, meta-analysis, weak instrumental variables, Mendelian randomization
1 Introduction
@@
Projecting long-term treatment effects from short-term metric movements is a ubiquitous problem in experimentation. In industrial settings, the problem is that businesses seek to optimize insensitive primary metrics (such as habitual usage, subscriber retention, and long-term revenue), but are unable or unwilling to measure treatment effects on these metrics precisely. [^1] For example, they may be interested in the effect of digital platform design on long-term user retention (Hohnhold et al., 2015), but they may be unwilling to run a sufficiently large experiment for a long time to measure treatment effects on long-term retention. To overcome this problem, they instead optimize for a suite of secondary metrics associated with those primary metrics, but which are more sensitive in terms of signal-to-noise and easier to measure in short-term experiments. Under certain assumptions, treatment effects on a “surrogate index” of multiple secondary metrics yields a precise estimate of the long-term treatment effect Athey et al. (2016); Prentice (1989).
@@
A class of statistical parameters that intuitively relates to the problem of constructing proxy metrics for a primary metric is the covariance matrix of true average treatment effects (ATEs) on primary and secondary metrics in previous experiments and functions thereof. For example, when constructing weighted indices of secondary outcomes, it is intuitive to consider the Ordinary (OLS) and Total Least Squares (TLS) regression of true ATEs on a primary outcome on true ATEs on the secondary outcomes in the scatterplot of true ATEs over available historical experiments, where by true ATE we mean the unobserved mean on the population, as compared to the estimated ATE actually observed on the experimental sample.
@@
How these statistical parameters actually connect to the question of proxy metrics or surrogates is the first question we investigate in this paper. In Section 3, we demonstrate that statistical features of the covariance matrix of true ATEs have causal interpretations under different causal models. In particular, in these models, they can support inference on effects of novel treatments on long-term outcomes based on short-term observations.
@@
The second question we study is that of estimating the covariance matrix of the true ATEs when the signal-to-noise ratio is small in each experiment, as is often the case in digital experimentation. The statistical challenges are analogous to the many weak instrumental variables (IVs) setting. One question of interest in the weak instrument literature is whether letting the number of instruments diverge while holding their strength fixed yields consistent estimates. Similarly, we demonstrate that by letting the number of experiments diverge, although the signal-to-noise ratio in each experiment remains bounded, we can consistently estimate the covariance matrix of true treatment effects. Each of the three methods we study actually has a weak-instrument estimator counterpart.
@@
Importantly, our results demonstrate that we can reliably estimate the covariance matrix of true treatment effects as a parameter at the meta-analytic level, that is, from experiment-level aggregates. In doing so, we contribute to an emerging literature on meta-analytic approaches to surrogacy Elliott et al. (2015); Cunningham and Kim (2020); Tripuraneni et al. (2023), which are well-suited to the large number of experiments conducted on modern online experimentation platforms. The data volume of historical experiments in online platforms is so massive that working at the unit level is often computationally prohibitive. Our methods are therefore also operationally, not just statistically, viable even in such large-scale experimentation settings.
@@
  • In Section 4, we present weak-IV-inspired estimators of the covariance matrix of true treatment effects, and of the OLS and TLS estimands in the true-ATEs scatterplot.
  • In Section 5, we conduct a simulation study to illustrate the performance of our proposed estimators and to provide visual intuition on the mechanics of some of our estimators.
  • Lastly, in Section LABEL:sec:linear_models, we describe the real-world application of our methods to derive a linear surrogate index for experimentation at Netflix
2 Statistical setup and notation
2.1 Data
We observe $N$ unit-level quadruples $O=(T,A,S,Y)$ where $T\in\{1,\ldots,K\}$ is the experiment (or test) index, $A\in\{0,1\}$ is the treatment arm index, $S$ is a $(G-1)$-dimensional vector of secondary metrics, and $Y$ is the primary metric of interest. We observe $N=Kn$ units divided between $K$ experiments, each of which has two arms with $n$ units. [^2] While our results extend to the more realistic setting of multiple treatment arms of varying size per experiment, we focus on the two-arm, constant sample size case for ease of exposition. Units in different experiments may be drawn from different superpopulations, but are assigned uniformly completely at random to cells. The observations $O_{1},\dots,O_{N}\sim O$ are each identically distributed. They may not be independent due to the restriction of having exactly $n$ units in each test. Conditional on $T_{1},\ldots,T_{N}$, any two observations are nonetheless independent.
2.2 Treatment-effect Covariance Matrix
For any $t=1,\ldots,K$, let $\tau_{S}(t)=E[S\mid A=1,T=t]-E[S\mid A=0,T=t]$ and $\tau_{Y}(t)=E[Y\mid A=1,T=t]-E[Y\mid A=0,T=t]$ be the true ATEs on the vector $S$ and on the scalar $Y$, respectively. Let
$\Lambda_{K}=\begin{bmatrix}\Lambda_{YY,K}&\Lambda_{SY,K}^{T}\\ \Lambda_{SY,K}&\Lambda_{SS,K}\end{bmatrix}$ (3)
@@ 2.2 Treatment-effect Covariance Matrix
$\Lambda_{YY,K}=$ $\frac{1}{K}\sum_{t=1}^{K}\tau_{Y}(t)^{2}-\left(\frac{1}{K}\sum_{t=1}^{K}\tau_{Y}(t)\right)^{2},$ (4)
$\Lambda_{SY,K}=$ $\frac{1}{K}\sum_{t=1}^{K}\tau_{S}(t)\tau_{Y}(t)-\left(\frac{1}{K}\sum_{t=1}^{K}\tau_{S}(t)\right)\left(\frac{1}{K}\sum_{t=1}^{K}\tau_{Y}(t)\right),$ (5)
$\Lambda_{SS,K}=$ $\frac{1}{K}\sum_{t=1}^{K}\tau_{S}(t)\tau_{S}(t)^{\top}-\left(\frac{1}{K}\sum_{t=1}^{K}\tau_{S}(t)\right)\left(\frac{1}{K}\sum_{t=1}^{K}\tau_{S}(t)\right)^{\top}.$ (6)
@@ 2.3 Statistical Parameters
$\theta_{1}(\Lambda_{K})=\Lambda_{SS,K}^{-1}\Lambda_{SY,K}.$ (7)
This is the OLS in the scatterplot of the $K$ true treatment effects. Our second parameter is
$\theta_{2,\Psi}(\Lambda_{K})=-\gamma_{\Psi,S}(\Lambda_{K})/\gamma_{\Psi,Y}(\Lambda_{K}),$ (8)
where $[\gamma_{\Psi,Y}(\Lambda_{K}),\gamma_{\Psi,S}(\Lambda_{K})^{T}]^{T}$ is a generalized eigenvector solving $(\Lambda_{K}-\kappa\Psi)\gamma=0$ for the smallest possible $\kappa\geq 0$ for which a solution exists, where $\Psi$ is a given positive definite matrix. The parameter $\theta_{2,\Psi}(\Lambda_{K})$ is the TLS on the $\Psi^{-1/2}$-transformed scatterplot of true treatment effects.
We emphasize that this OLS and TLS are defined in terms of only $K$ points in a $G$-dimensional scatterplot. The location of these points themselves (the true ATEs) are, nonetheless, unknown population quantities. Therefore, $\theta_{1}(\Lambda_{K})$ is a population quantity, that is, it is a function of the unknown population distribution of the data $O$ (equivalently, of infinite draws of $O$).
3 Causal Models and Relationship to Proxies
We now consider various causal models and see how the statistical parameters $\theta_{1}(\Lambda_{K})$ and $\theta_{2}(\Lambda_{K})$ relate to parameters in these causal models, as well as support the construction of proxy metrics.
3.1 Linear Structural Model without Direct Effects
@@
$S$ $=\mu_{S}(T)+\pi_{S}(S)A+\gamma U+\eta$ (9)
$Y$ $=\mu_{Y}(T)+S\beta+\delta U+\epsilon.$ (10)
@@
$\beta=\theta_{1}(\Lambda_{K})=\theta_{2,\Psi}(\Lambda_{K}),$ (11)
for any positive definite $\Psi$. The equality of the two statistical parameters can be understood as follows. Under the DGP induced by the above structural model, in which $S$ fully mediates the effect of $A$ on $Y$, $\tau_{Y}(t)=\tau_{S}(t)\beta$ for every $t\in[K]$, which implies that $[1,-\beta]^{T}$ is an eigenvector associated with eigenvalue 0, which must be the smallest as $\Lambda_{K}$ is positive semi-definite. Therefore, the TLS estimand is also the OLS estimand.
@@
Now consider a new experiment $T=K+1$, where we also randomize units equiprobably between treatment ($A=1$) and control ($A=0$). Denoting $S(t,a)$, $Y(t,a)$ the potential outcomes generated by the structural model (that is, the random variables obtained by setting values $T=t$ and $A=a$ and sampling $U$ and the error terms in the above equations), the ATE on $Y$ is related to the difference in arm-specific means on $S$ as follows:
$E[Y(K+1,1)-Y(K+1,0)]$ (12)
$=E[S(K+1,1)-S(K+1,0)]\beta$ (13)
$=(E[S\mid T=K+1,A=1]-E[S\mid T=K+1,A=0])\theta_{i}(\Lambda_{K}),$ (14)
@@
3.2 Linear Structural Model with INSIDE-consistent Direct Effects
@@
$S$ $=\mu_{S}(T)+\pi_{S}(T)A+\gamma U+\eta$ (15)
$Y$ $=\mu_{Y}(T)+\pi_{Y}(T)A+S\beta+\delta U+\epsilon.$ (16)
It is generally impossible to disentangle the direct ($\pi_{Y}(T)$) and indirect/mediated ($\pi_{S}(T)\beta$) effects. To overcome this, we make the INstrument Strength Independent of Direct Effect (INSIDE) assumption, originating from the Mendelian randomization literature Burgess and Thompson (2017). In particular, we impose that the vector $\bm{\pi}_{Y}=[\pi_{Y}(1),\ldots,\pi_{Y}(K)]^{\top}$ of direct effects is orthogonal to the columns of the matrix $\bm{\Pi}_{S}=[\pi_{S}(1)^{\top},\ldots,\pi_{S}(K)^{T}]^{\top}$ of first-stage effects on $S$, that is, $\bm{\pi}_{Y}^{T}\bm{\Pi}_{S}=0$.
@@
$\Lambda_{K}=\begin{bmatrix}\beta^{T}\Lambda_{SS,K}\beta&(\Lambda_{SS,K}\beta)^{\top}\\ \Lambda_{SS,K}\beta&\Lambda_{SS,K}\end{bmatrix}+\begin{bmatrix}\frac{\pi_{Y}^{\top}\pi_{Y}}{K}-\frac{(\pi_{Y}^{\top}\bm{1})^{2}}{K^{2}}&0\\ 0&0,\end{bmatrix}$ (21)
@@
As can be seen from the above matrix expression, $\theta_{1}(\Lambda_{K})=\beta$ identifies the structural parameter $\beta$. However, it is no longer the case in general that $[1,-\beta]^{\top}$ is an eigenvector of $\Lambda_{K}$ and therefore the TLS estimand $\theta_{2,\Psi}(\Lambda_{K})$ diverges in general from $\beta$.
@@
In the presence of direct effects, it is no longer the case that ATEs on $Y$ are given by ATEs on $S$ times $\beta$, as $S$ no longer mediates the effect of the treatment fully. Still, $\beta$, which coincides with the OLS estimand $\theta_{1}(\Lambda_{K})$, has a causal interpretation under the linear structural model. Let us introduce the potential outcomes $S(t,a)$ and $Y(t,a,s)$ generated by the structural model (that is, again, the random variables obtained by setting $T=t,A=a,S=s$ in the above equations and sampling independently $U$ and the error terms). It holds that the natural indirect effect on $Y$ through $S$ is given by
$E[Y(K+1,a,S(K+1,1))-Y(K+1,a,S(K+1,0)]$ (22)
$=E[S(K+1,1)-S(K+1,0)]\beta$ (23)
$=(E[S\mid T=K+1,A=1]-E[S\mid T=K+1,A=0])\theta_{1}(\Lambda_{K}),$ (24)

### v2

@@ Abstract.
When primary objectives are insensitive or delayed, experimenters may instead focus on proxy metrics derived from secondary outcomes. For example, technology companies often infer the long-term impacts of product interventions from their effects on short-term user engagement signals. We consider the meta-analysis of many historical experiments to learn the covariance of treatment effects on these outcomes, which can support the construction of such proxies. Even when experiments are plentiful, if treatment effects are weak, the covariance of estimated treatment effects across experiments can be highly biased. We overcome this with techniques inspired by weak instrumental variable analysis. We show that Limited Information Maximum Likelihood (LIML) learns a parameter equivalent to fitting total least squares to a transformation of the scatterplot of treatment effects, and that Jackknife Instrumental Variables Estimation (JIVE) learns another parameter computable from the average of Jackknifed covariance matrices across experiments. We also present a total covariance estimator for the latter estimand under homoskedasticity, which is equivalent to a $k$-class estimator. We show how these parameters can be used to construct unbiased proxy metrics under various structural models. Lastly, we discuss the real-world application of our methods at Netflix.
surrogate outcomes, meta-analysis, weak instrumental variables
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3672034isbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Learning linear modelsccs: Mathematics of computing Multivariate statistics
1. Introduction
@@
Projecting long-term treatment effects from short-term metrics is a ubiquitous problem in experimentation. For example, technology companies seek to optimize insensitive primary metrics (such as habitual usage, subscriber retention, and long-term revenue), but are unable or unwilling to measure treatment effects on these metrics precisely. [^1] For example, they may be interested in the effect of digital platform design on long-term user retention (hohnhold2015focusing, ), but unwilling to run a sufficiently large experiment for a long time to measure treatment effects on long-term retention. To address this issue, they optimize a suite of secondary metrics that are associated with the primary metrics, but more sensitive in terms of signal-to-noise and easier to measure in short-term experiments. Under certain assumptions, the treatment effect on a “surrogate index” of multiple secondary metrics yields a precise estimate of the long-term treatment effect (athey2016estimating, ; prentice1989surrogate, ).
@@
A class of statistical parameters that intuitively relates to the problem of constructing proxy metrics for a primary metric is the covariance matrix of true average treatment effects (ATEs) on primary and secondary metrics in previous experiments, and functions thereof. For example, when constructing weighted indices of secondary outcomes, it is intuitive to consider the Ordinary (OLS) and Total Least Squares (TLS) regression of true ATEs on a primary outcome on true ATEs on the secondary outcomes in the scatterplot of true ATEs over available historical experiments. Here, by true ATE, we mean the unobserved mean on the population, in contrast to the estimated ATE actually observed on the experimental sample.
@@
How these statistical parameters actually connect to the question of proxy metrics or surrogates is the first question we investigate in this paper. In Section 3, we demonstrate that statistical features of the covariance matrix of true ATEs have causal interpretations under different causal models. For instance, under these models, they can support inference on effects of novel treatments on long-term outcomes based on short-term observations.
@@
The second question we study is the estimation of the covariance matrix of the true ATEs when the signal-to-noise ratio is small in each experiment, as is often the case in digital experimentation (peysakhovich2018learning, ; deng2013improving, ; xie2016improving, ). The statistical challenges are analogous to the many weak instrumental variables (IVs) setting. One question of interest in the weak instrument literature is whether allowing the number of instruments to diverge while holding their strength fixed yields consistent estimates (mikusheva2022inference, ). Analogously, we show that increasing the number of experiments, even if each maintains a low signal-to-noise ratio, enables consistent estimation of the covariance matrix of true treatment effects. In fact, each of the three methods we study has a weak-instrument estimator counterpart.
@@
Our results demonstrate that we can reliably estimate the covariance matrix of true treatment effects as a parameter at the meta-analytic level, that is, from experiment-level aggregates. In this way, we contribute to an emerging literature on meta-analytic approaches to surrogacy (elliott2015surrogacy, ; cunningham2020interpreting, ; tripuraneni2023choosing, ), which is particularly relevant given the large number of experiments conducted on modern online experimentation platforms. The sheer volume of historical data from these platforms often makes unit-level analysis computationally challenging, if not prohibitive (wong2019efficient, ). Therefore, our methods are not only statistically robust, but also operationally feasible for large-scale experimentation.
@@
  • In Section 4, we present weak-IV-inspired estimators of the covariance matrix of true treatment effects and the OLS and TLS estimands in the scatterplot of true ATEs.
  • In Section 5, we conduct a simulation study to illustrate the performance of our proposed estimators and provide visual intuition on the mechanics of some of our estimators.
  • Lastly, in Section 6, we describe the real-world application of our methods to derive a linear surrogate index for experimentation at Netlix.
2. Statistical setup and notation
2.1. Data
We observe $N$ unit-level quadruples $O=(T,A,S,Y)$ where $T\in\{1,\ldots,K\}$ is the experiment index, $A\in\{0,1\}$ is the treatment arm index, $S$ is a $(G-1)$-dimensional vector of secondary metrics, and $Y$ is the primary metric of interest. We observe $N=Kn$ units divided between $K$ experiments, each of which has two arms with $n$ units. [^2] While our results extend to the more realistic setting of multiple treatment arms of varying size per experiment, we focus on the two-arm, constant sample size case for ease of exposition. Units in different experiments may be drawn from different superpopulations, but are assigned uniformly completely at random to cells. The observations $O_{1},\dots,O_{N}\sim O$ are each identically distributed. They may not be independent due to the restriction of having exactly $n$ units in each experiment. Conditional on $T_{1},\ldots,T_{N}$ and $A_{1},\ldots,A_{N}$, any two observations are nonetheless independent.
2.2. Treatment-effect Covariance Matrix
For any $t=1,\ldots,K$, let $\tau_{S}(t)=E[S\mid A=1,T=t]-E[S\mid A=0,T=t]$ and $\tau_{Y}(t)=E[Y\mid A=1,T=t]-E[Y\mid A=0,T=t]$ be the true ATEs in experiment $t$ on the vector $S$ and on the scalar $Y$, respectively. Let
(1) $\Lambda_{K}=\begin{bmatrix}\Lambda_{YY,K}&\Lambda_{SY,K}^{T}\\ \Lambda_{SY,K}&\Lambda_{SS,K}\end{bmatrix}$
@@ 2.2. Treatment-effect Covariance Matrix
(2) $\Lambda_{YY,K}=$ $\frac{1}{K}\sum_{t=1}^{K}\tau_{Y}(t)^{2}-\left(\frac{1}{K}\sum_{t=1}^{K}\tau_{Y}(t)\right)^{2},$
(3) $\Lambda_{SY,K}=$ $\frac{1}{K}\sum_{t=1}^{K}\tau_{S}(t)\tau_{Y}(t)-\left(\frac{1}{K}\sum_{t=1}^{K}\tau_{S}(t)\right)\left(\frac{1}{K}\sum_{t=1}^{K}\tau_{Y}(t)\right),$
(4) $\Lambda_{SS,K}=$ $\frac{1}{K}\sum_{t=1}^{K}\tau_{S}(t)\tau_{S}(t)^{\top}-\left(\frac{1}{K}\sum_{t=1}^{K}\tau_{S}(t)\right)\left(\frac{1}{K}\sum_{t=1}^{K}\tau_{S}(t)\right)^{\top}.$
@@ 2.3. Statistical Parameters
(5) $\theta_{1}(\Lambda_{K})=\Lambda_{SS,K}^{-1}\Lambda_{SY,K}.$
This is the OLS in the scatterplot of the $K$ true treatment effects. While we discuss the exact causal interpretation of this parameter (which requires structural assumptions) in the next section, intuitively the OLS measures the statistical relationship between the surrogate metrics and the long-term metric (tripuraneni2023choosing, ).
Our second parameter is more complex:
(6) $\theta_{2,\Psi}(\Lambda_{K})=-\gamma_{\Psi,S}(\Lambda_{K})/\gamma_{\Psi,Y}(\Lambda_{K}),$
where $[\gamma_{\Psi,Y}(\Lambda_{K}),\gamma_{\Psi,S}(\Lambda_{K})^{T}]^{T}$ is a generalized eigenvector solving $(\Lambda_{K}-\kappa\Psi)\gamma=0$ for the smallest possible $\kappa\geq 0$ for which a solution exists, where $\Psi$ is a given positive definite matrix. The parameter $\theta_{2,\Psi}(\Lambda_{K})$ is the TLS on the $\Psi^{-1/2}$-transformed scatterplot of true treatment effects. While this parameter is less intuitive, we show below that it coincides with $\theta_{1}(\Lambda_{K})$ under certain structural assumptions.
We emphasize that these OLS and TLS parameters are defined in terms of only $K$ points in a $G$-dimensional scatterplot. The location of these points themselves (the true ATEs) are, nonetheless, unknown population quantities. Therefore, $\theta_{1}(\Lambda_{K})$ is a population quantity, meaning it is a function of the unknown population distribution of the data $O$ (that is, of infinite draws of $O$).
3. Causal Models and Relationship to Proxies
We now consider various causal models and examine how the statistical parameters $\theta_{1}(\Lambda_{K})$ and $\theta_{2}(\Lambda_{K})$ relate to causal parameters in these models. Equivalently, this section derives the structural assumptions under which these statistical parameters can be given a causal interpretation in terms of the effect of $S$ on $Y$ and thereby used to construct proxy metrics for $Y$.
3.1. Linear Structural Model without Direct Effects
@@
(7) $S$ $=\mu_{S}(T)+\pi_{S}(T)A+\gamma U+\eta$
(8) $Y$ $=\mu_{Y}(T)+S\beta+\delta U+\epsilon.$
Here, $\pi_{S}(T)$ is the $(G-1)\times 1$ vector of first-stage effects in experiment $T$ (that is, the ATEs on the $G-1$ secondary metrics $S$ of the intervention trialed in experiment $T$), $\mu_{S}(T)$ is a $(G-1)\times 1$ vector of per-experiment fixed effects on $S$, and $\mu_{Y}(T)$ is a scalar per-experiment fixed effect on $Y$.
@@
(9) $\beta=\theta_{1}(\Lambda_{K})=\theta_{2,\Psi}(\Lambda_{K})$
for any positive definite $\Psi$. The equality of the two statistical parameters can be understood as follows. Under the DGP induced by the above structural model, in which $S$ fully mediates the effect of $A$ on $Y$, $\tau_{Y}(t)=\tau_{S}(t)\beta$ for every $t\in[K]$. This implies that $[1,-\beta]^{T}$ is an eigenvector associated with eigenvalue 0, which must be the smallest as $\Lambda_{K}$ is positive semi-definite. Therefore, the TLS estimand is also the OLS estimand.
@@
Now consider a new experiment $T=K+1$, in which we also randomize units equiprobably between treatment ($A=1$) and control ($A=0$). Denoting $S(t,a)$, $Y(t,a)$ the potential outcomes generated by the structural model (that is, the random variables obtained by setting values $T=t$ and $A=a$ and sampling $U$ and the error terms in the above equations), the ATE on $Y$ is related to the difference in arm-specific means on $S$ as follows:
(10) $E[Y(K+1,1)-Y(K+1,0)]$
(11) $=E[S(K+1,1)-S(K+1,0)]\beta$
(12) $=(E[S\mid T=K+1,A=1]-E[S\mid T=K+1,A=0])\theta_{i}(\Lambda_{K}),$
@@
3.2. Linear Structural Model with INSIDE-consistent Direct Effects
@@
(13) $S$ $=\mu_{S}(T)+\pi_{S}(T)A+\gamma U+\eta$
(14) $Y$ $=\mu_{Y}(T)+\pi_{Y}(T)A+S\beta+\delta U+\epsilon.$
It is generally impossible to disentangle the direct ($\pi_{Y}(T)$) and indirect/mediated ($\pi_{S}(T)\beta$) effects. Still, it is possible to estimate meaningful causal parameters under the assumption of INstrument Strength Independent of Direct Effect (INSIDE) from the Mendelian randomization literature (burgess2017interpreting, ), which states that the first-stage effects of $A$ on $S$ are independent of the direct effects of $A$ on $Y$. In particular, INSIDE requires that the vector $\bm{\pi}_{Y}=[\pi_{Y}(1),\ldots,\pi_{Y}(K)]^{\top}$ of direct effects is orthogonal to the columns of the matrix $\bm{\Pi}_{S}=[\pi_{S}(1)^{\top},\ldots,\pi_{S}(K)^{T}]^{\top}$ of first-stage effects on $S$, that is, $\bm{\pi}_{Y}^{T}\bm{\Pi}_{S}=0$.
@@
(15) $\Lambda_{K}=\begin{bmatrix}\beta^{T}\Lambda_{SS,K}\beta&(\Lambda_{SS,K}\beta)^{\top}\\ \Lambda_{SS,K}\beta&\Lambda_{SS,K}\end{bmatrix}+\begin{bmatrix}\frac{\pi_{Y}^{\top}\pi_{Y}}{K}-\frac{(\pi_{Y}^{\top}\bm{1})^{2}}{K^{2}}&0\\ 0&0,\end{bmatrix}$
@@
As can be seen from the above matrix expression, $\theta_{1}(\Lambda_{K})=\beta$ identifies the structural parameter $\beta$. However, with direct effects, it is no longer the case in general that $[1,-\beta]^{\top}$ is an eigenvector of $\Lambda_{K}$ and therefore the TLS estimand $\theta_{2,\Psi}(\Lambda_{K})$ diverges in general from $\beta$.
@@
In the presence of direct effects, it is no longer the case that ATEs on $Y$ are given by ATEs on $S$ times $\beta$, as $S$ no longer mediates the effect of the treatment fully. Still, $\beta$, which coincides with the OLS estimand $\theta_{1}(\Lambda_{K})$, has a causal interpretation under the linear structural model. Let us introduce the potential outcomes $S(t,a)$ and $Y(t,a,s)$ generated by the structural model (that is, the random variables obtained by setting $T=t,A=a,S=s$ in the above equations and independently sampling $U$ and the error terms). It holds that the natural indirect effect on $Y$ through $S$ is given by
(16) $E[Y(K+1,a,S(K+1,1))-Y(K+1,a,S(K+1,0)]$
(17) $=E[S(K+1,1)-S(K+1,0)]\beta$
(18) $=(E[S\mid T=K+1,A=1]-E[S\mid T=K+1,A=0])\theta_{1}(\Lambda_{K}),$